OpenStudy (anonymous):

Traffic signs are regulated by the manual on uniform traffic control devices (MUTCD)the perimeter of a rectangular traffic sign is 76 inches. Also, it's length is 12 inches longer than its width. Find the dimensions of this sign. The width of the sign is___in. The length of the sign is___in.

OpenStudy (anonymous):

im lost pl help me

OpenStudy (anonymous):

length is 12 inches longer than width therefore l=12+w perimeter = 2(l+w) =2(12+w+w)=76 ;)

OpenStudy (anonymous):

19in w6

OpenStudy (anonymous):

i think when you add 12 in 6 , its 18 ;)

OpenStudy (anonymous):

ok18im w6

OpenStudy (anonymous):

cool ;)

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

13in 25w

OpenStudy (anonymous):

w=13 , and length = 25

OpenStudy (anonymous):

i got a red x on it